if sinx+sin2x=1 then explain how to solve


cos2x +cos4x=?


Grade:12

3 Answers

Godfrey Classic Prince
12 years ago

Dear Abhijat Pandey,

cos2x + cos4x =12

I will give my explaination in the other answer shortly ..!

 

Thank You.

Please approve my answer if you liked it by clicking on "Yes" given below...!!Smile

Bajrangi Baahubali
8 years ago
 
cos^2x+cos^4x=? …...(1)
 
sinx+sin^2x =1
since, sin^2x+cos^2x=1
therefore, sin^2x=1-cos^2x
hence, sinx+sin^2x becomes sinx+1-cos^2x=1
which concludes sinx=cos^2x  …..(2)
similarly, sin^2=cos^4x (by Squaring (2))
on substituting
(2) in (1)
we get sinx+sin^2=cos^2x+cos^4x=1
 
 
Akshat
6 years ago
Bla bla bla bla bla bla bla blaBla bla bla bla blaBla bla blaBla bla Bla BlB0+1Bla bla bla bla bla blaBla bla bla bla blaBla bla blaBla blaBla blaBl B0